Identify the slope and y-intercept. graph the equation. y=-1/3x-3 *please note that the "x" is a variable.*
Really struggling with this question.

Answer by oscargut(2103)


Identify the slope and y-intercept. graph the equation. y=-1/3x-3 *please note that the "x" is a variable.*

Answer:
Slope = -1/3
y-intercept = (0,-3)
You can ask me more at: mthman@gmail.com
Thanks